OpenStudy (anonymous):

Prove that 1-2sin^2( pi/4 - x/2) = sin x

OpenStudy (valpey):

\[A:1 = \cos^2+\sin^2\]\[B:\cos^2x-\sin^2x = \cos(2x)\]\[C:\cos{\frac{\pi}{2}-x}=\sin x\]

OpenStudy (anonymous):

?? .